Marie can paint a room in 9 hours. If Ana helps, it takes them 7 hours. Without help, how long would it take Ana to complete this job?
Answer by Boreal(15235)

In 7 hours Marie paints the equivalent of 7/9 of the room.
So Ana paints 2/9 the equivalent in 7 hours.
(2/9)/7=1/x
7=(2/9)*x
x=31.5 hours for Ana.
Can check (x/9)+(x/31.5)=1
31.5x+9x=283.5, multiplying by 283.5
40.5x=283.5
x=7
It will take Ana 31.5 hours alone.